Circular No. 7314

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999fz IN UGC 8164 T. Puckett, A. Langoussis, and D. George, Mountain Town, GA, re-

port their discovery of an apparent supernova (mag 15.2) on an unfiltered CCD frame (limiting mag 20.1) taken with the Puckett Observatory 0.60-m automated supernova patrol telescope on Nov. 18.40 UT. SN 1999fz is located at α = 13^{h}02^{m}16^{s}.34, δ = +78^{o}32′29′′.4 (equinox 2000.0), which is 2′′.0 west and 1′′.3 south of the center of UGC 8164. SN 1999fz was also present on unfiltered CCD frames taken on Nov. 21.98, but it does not appear on a Palomar Sky survey image (limiting mag 20.0) taken on 1952 Oct. 22.

HD 209458 R. Rebolo, M. R. Zapatero Osorio, H. J. Deeg, and J. Licandro, Insti-

tuto de Astrofísica de Canarias; and J. Martí, University of Jaen, report: “CCD photometry of the star HD 209458 obtained at the 1.5-m Spanish Telescope at Calar Alto reveals a transit egress of the planet reported on IAUC 7307, with a depth of 0.023 ± 0.005 mag, on Nov. 18.87 UT. Ingress started before observations, but assuming an orbital period of 3.523 days, the transit lasts 3.0 ± 0.1 hr. Best fits give an orbital inclination of 87^{o}.0 ± 0^{o}.5 and a planet radius 1.34 ± 0.07 that of Jupiter.”

R. A. Street, T. A. Lister, A. Cameron, and K. Horne, St. Andrews University, report that 640-nm narrowband CCD photometry of HD 209458 with the 0.9-m Gregory telescope in St. Andrews confirms that a planetary transit with depth of 0.019 ± 0.003 mag occurred from Nov. 18.760 ± 0.021 to 18.903 ± 0.031 UT. Z. Ioannou, C. Hellier, R. D. Jeffries, and T. Naylor, Keele University, report that V -band photometry with the 0.6-m telescope at Keele Observatory are consistent with an egress as reported above, but the increasing airmass makes the data from this time onwards unreliable. For further details, see http://star-www.st-

and.ac.uk/~ras11/indexPS.html.”

2S 1417–624

Corrigendum. On IAUC 7313, line 7, for 56.99490(9) Hz. read 56.99490(9) mHz.

1999 November 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green